OpenStudy (anonymous):

how do i graph y= log(base 5)x??????????

OpenStudy (campbell_st):

the same way that you graph \[\log_{10}(x).... or ..... \log_{e}(x)\]  if you wanted a table of values you could use change of base to something your calculator can handle \[\log_{5}(x) = \frac{\ln(x)}{\ln(5)}\] then use a table of values like x: 0.1 0.2 0.5 1 2 5 10 y to find the y values substitute the x values into the change of base information then plot the ordered pairs